1. Expected Return If a company's current stock price is $26.10 and it is likely to pay a $1.85 dividend next year. Since analysts estimate the company will have a 15% growth rate, what is its expected return?

  • 2.13%

  • 7.09%

  • 15.00%

  • 22.09%

2. Average Return The past five monthly returns for K and Company are 4.85 percent, 4.73 percent, -1.45 percent, 3.85 percent, and 7.85 percent. What is the average monthly return?

  • 3.966%

  • 1.894%

  • 4.546%

  • 1.653%

3. Average Return The past five monthly returns for K and Company are 3.85 percent, 4.02 percent, -1.35 percent, -.85 percent, and 8.60 percent. What is the average monthly return?

  • 3.734%

  • 1.556%

  • 2.854%

  • 1.189%
Answer #1

Ans 1) 22.09%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
26.10 = 1.85 / (Ke- 15%)
Ke - 15% = 1.85 / 26.10
Ke - 15% = 7.09%
Ke = 22.09%

Ans 2) 3.966%

Period RETURN
1 4.85
2 4.73
3 -1.45
4 3.85
5 7.85
Total 19.83
Average Return = Total Return / NO of Periods
19.83 / 5
3.966

Ans 3) 2.854%

Period RETURN
1 3.85
2 4.02
3 -1.35
4 -0.85
5 8.6
Total 14.27
Average Return = Total Return / NO of Periods
14.27 / 5
2.854